# Forum for Information Retrieval Evaluation, December 12-15, 2024, India

### v1

@@ 1 Introduction
Code-mixing, where elements from multiple languages are blended within a single sentence, is a natural and widespread phenomenon in multilingual societies [1, 2, 3]. It is particularly prevalent in India, a country with a rich linguistic diversity where speakers often switch between languages depending on context, audience, and medium of communication [4, 5]. With the rapid rise of online social networking, this practice has become increasingly common in digital conversations, where users frequently combine their native languages with others, often using foreign scripts [6, 7].
One notable trend in India is the use of the Roman script to communicate in native languages on social media platforms [8, 9]. This practice is especially common among migrant communities who form online groups to share information and experiences relevant to their unique circumstances [10, 11]. For instance, Bengali speakers from West Bengal who have migrated to urban centers like Delhi or Bangalore often establish groups such as "Bengali in Delhi" on platforms like Facebook and WhatsApp. These groups serve as vital hubs for exchanging advice on a wide range of local issues, from housing and employment to navigating new social environments.
The COVID-19 pandemic highlighted the importance of these online communities as critical sources of information [12, 13]. During this period, these groups became essential for sharing experiences, seeking support, and keeping up with the frequently changing government guidelines. However, the informal and often colloquial nature of the language used in these code-mixed conversations, typically transliterated into Roman script, presents significant challenges for information retrieval. The lack of standardization, combined with the blending of languages, makes it difficult to identify and extract relevant answers, especially for those who might seek similar information at a later time [14, 15].
This paper addresses the challenge of extracting relevant information from code-mixed digital conversations, with a specific focus on Roman transliterated Bengali mixed with English. While code-mixing is a well-recognized phenomenon in natural language processing (NLP), the unique characteristics of transliterated text—such as variations in spelling, grammar, and syntax—complicate the task of effective information retrieval [16, 17]. To tackle this issue, we have developed a mechanism that identifies the most relevant answers from these complex, multilingual discussions.
@@
We leverage GPT-3.5 Turbo [18, 19] by employing carefully designed prompts that guide the model to evaluate the relevance of documents with respect to a given query. This involves not only the semantic understanding capabilities of GPT-3.5 Turbo but also the strategic use of the sequential nature of documents. Often, documents are part of a series or a conversation where the relevance to a query can be influenced by preceding or succeeding documents. By acknowledging this sequence, we can better capture contextual relationships that might be missed if documents were considered in isolation.
@@ 2 Related Work
Code-mixing and transliteration have gained increasing attention in the field of natural language processing (NLP), especially as global communication becomes more digital and multilingual [20, 21, 22, 23]. This section reviews key studies related to code-mixing, information retrieval from code-mixed text, and the challenges of processing Roman transliterated languages, particularly in the context of Indian languages. Code-mixing, where speakers blend elements from multiple languages within a single utterance, is a common linguistic phenomenon in multilingual societies [22, 24, 25]. Early studies on code-mixing focused primarily on sociolinguistic aspects, examining how and why speakers switch languages within conversations [20, 21, 22, 26, 27]. However, with the advent of digital communication, researchers have increasingly turned their attention to computational methods for processing and understanding code-mixed text [28, 29, 30].
@@
Several studies have explored various NLP tasks, such as part-of-speech tagging, language identification, and sentiment analysis, in code-mixed settings [31, 32]. [33, 34] provided one of the earliest comprehensive analyses of code-mixed text, highlighting the unique challenges it poses for traditional NLP pipelines, such as non-standard spelling, syntax variations, and the blending of multiple languages within a single text. More recent work by [35, 36, 37] introduced a code-mixed dataset, spanning multiple Indian languages, which has become a benchmark for evaluating NLP models in this domain.
@@
Information retrieval (IR) in code-mixed settings is relatively underexplored compared to other NLP tasks [38, 39, 40]. However, the need for effective IR systems that can handle multilingual and code-mixed queries has become increasingly important, particularly in the context of digital information exchange on social media platforms. [41, 40] investigated the problem of query-focused summarization in code-mixed social media data, emphasizing the complexity of extracting relevant information from noisy, informal text. Work by [42] addressed code-mixed question answering, where the goal is to identify correct responses from a mixed-language corpus. Their approach involved using translation models to standardize the text before applying traditional IR techniques, demonstrating that even simple translation-based methods can significantly improve performance. However, these methods often fail to capture the nuances of code-mixed language, such as cultural context and colloquial expressions.
@@
Roman script transliteration of Indian languages, commonly referred to as "Romanagari" for languages like Hindi, is a widespread practice in digital communication. Transliteration introduces additional challenges for NLP, as it often involves non-standard spellings and inconsistent usage. For instance, multiple transliterations may exist for the same word, depending on the speaker’s regional accent, literacy in the original script, or personal preference.
@@
Notable efforts in this area include the work by [43, 44], which explored transliteration normalization for Hindi-English code-mixed text. They developed algorithms to map Romanized text back to its original script, enabling more accurate processing by traditional NLP models. However, normalization remains a challenging task due to the inherent variability in transliterated text. In the context of Bengali, the Roman script transliteration is less standardized than for Hindi, leading to even greater variability in spelling and grammar. [45, 46] addressed this issue by creating a Roman Bengali dataset and proposed methods for transliteration normalization and language identification. Their work highlights the difficulties of processing Roman Bengali and the need for specialized approaches tailored to the characteristics of the language.
@@
Large Language Models (LLMs) [46, 47, 48, 49, 50, 51, 52] like GPT-3 have shown promise in various NLP tasks, including LI. Previous works have demonstrated the capability of GPT-3 in performing zero-shot and few-shot learning, making it a potentially powerful tool for LI in resource-constrained settings. However, the application of LLMs [53, 54, 55, 56, 57, 58, 59] to code-mixed and morphologically rich languages remains underexplored. Recent studies, have started to explore the use of transformers and pre-trained models for multilingual LI, but the effectiveness of these models in Bengali languages requires further investigation.
@@ 3 Dataset
This shared task consists of a single dataset for code mixed information retrieval. The corpus consists of 107900 documents in the training set and 20 queries in the training set. There are 30 queries in the testing set. The dataset is in roman transliterated bengali mixed with english language.
@@ 4 Task Definition
The task is to automatically determine the relevance of a query to a document within code-mixed data, specifically focusing on English and Roman transliterated Bengali.
@@ 5.1 Why Prompting?
  • Scalability and Customization: Prompting for information retrieval offers scalability and customization that traditional IR systems might lack. By designing prompts tailored to specific contexts or types of queries, LLMs can be dynamically adjusted to meet the needs of different retrieval tasks [64, 65]. This customization is particularly useful in handling domain-specific language or code-mixed scenarios, where standard IR systems might require extensive re-training or re-configuration.
@@ 5.2 Merging Prompt and Mathematical Model-Based Approaches
We used the GPT-3.5 Turbo model via prompting through the OpenAI API [^1] https://platform.openai.com/docs/models/gpt-3-5-turbo to solve the document retrieval task. We used the following prompt:
@@
For the five results reported, we ran the GPT model at different temperature values namely 0.5, 0,6, 0.7, 0.8, and 0.9. The diagram for GPT-3.5 Turbo is shown in Figure  1. The figure representing the methodology is shown in Figure  2.
@@ 6 Results
Table 1 presents the evaluation metrics for different submissions for the team named "TextTitans". The metrics used to assess the performance are MAP Score, ndcg Score, p@5 Score, and p@10 Score. Here’s what these results imply. MAP is a common metric in information retrieval that measures the precision of results across multiple queries. A higher MAP score indicates that relevant documents are consistently ranked higher across all queries. In the table, the MAP scores for the first four submissions are identical (0.701773), while the fifth submission slightly improves to 0.703734. This indicates that the fifth submission is marginally better in terms of ranking relevant results across multiple queries. The ndcg score measures the quality of the ranking based on the position of relevant documents. A higher ndcg score suggests that relevant documents are placed higher in the ranking. The scores are also very similar across submissions, with the first four submissions having an ndcg score of 0.797937, and the fifth submission showing a slight improvement to 0.799196. This suggests a minor improvement in ranking relevant documents for the fifth submission. p@5 measures how many of the top 5 ranked documents are relevant. A score of 1 would mean that all 5 of the top-ranked documents are relevant. All submissions have the same p@5 score of 0.793333, indicating that the top 5 results are equally accurate across all submissions. Precision@10 measures how many of the top 10 ranked documents are relevant. Like p@5, a higher score is better. Similar to p@5, all submissions have the same p@10 score of 0.766667, showing no variation in the top 10 results across the different submissions. The metrics are very consistent across all submissions, with only minor improvements in MAP and NDCG scores for the fifth submission. The fifth submission shows a slight improvement in ranking and retrieval performance, but the changes are minimal. The p@5 and p@10 scores indicate that the precision of the top 5 and top 10 results is identical across all submissions, suggesting that the models are performing similarly in identifying the most relevant documents. Overall, while there is a slight improvement in the last submission, the models generally perform similarly across all metrics.

### v3

@@ 1 Introduction
Code-mixing, where elements from multiple languages are blended within a single sentence, is a natural and widespread phenomenon in multilingual societies [1, 2]. Code mixing is a global phenomenon where speakers often switch between languages depending on context, audience, and medium of communication [3]. With the rapid rise of online social networking, this practice has become increasingly common in digital conversations, where users frequently combine their native languages with others, often using foreign scripts [4].
One notable trend in India is the use of the Roman script to communicate in native languages on social media platforms [5]. This practice is especially common among migrant communities who form online groups to share information and experiences relevant to their unique circumstances [6]. For instance, Bengali speakers from West Bengal who have migrated to urban centers like Delhi or Bangalore often establish groups such as "Bengali in Delhi" on platforms like Facebook and WhatsApp. These groups serve as vital hubs for exchanging advice on a wide range of local issues, from housing and employment to navigating new social environments.
The COVID-19 pandemic highlighted the importance of these online communities as critical sources of information [7]. During this period, these groups became essential for sharing experiences, seeking support, and keeping up with the frequently changing government guidelines. However, the informal and often colloquial nature of the language used in these code-mixed conversations, typically transliterated into Roman script, presents significant challenges for information retrieval. The lack of standardization, combined with the blending of languages, makes it difficult to identify and extract relevant answers, especially for those who might seek similar information at a later time [8].
This paper addresses the challenge of extracting relevant information from code-mixed digital conversations, with a specific focus on Roman transliterated Bengali mixed with English. While code-mixing is a well-recognized phenomenon in natural language processing (NLP), the unique characteristics of transliterated text—such as variations in spelling, grammar, and syntax—complicate the task of effective information retrieval [9]. To tackle this issue, we have developed a mechanism that identifies the most relevant answers from these complex, multilingual discussions.
@@
We leverage GPT-3.5 Turbo [10] by employing carefully designed prompts that guide the model to evaluate the relevance of documents with respect to a given query. This involves not only the semantic understanding capabilities of GPT-3.5 Turbo but also the strategic use of the sequential nature of documents. Often, documents are part of a series or a conversation where the relevance to a query can be influenced by preceding or succeeding documents. By acknowledging this sequence, we can better capture contextual relationships that might be missed if documents were considered in isolation.
@@ 2 Related Work
Code-mixing and transliteration have gained increasing attention in the field of natural language processing (NLP), especially as global communication becomes more digital and multilingual [11, 12, 13]. This section reviews key studies related to code-mixing, information retrieval from code-mixed text, and the challenges of processing Roman transliterated languages, particularly in the context of Indian languages. Code-mixing, where speakers blend elements from multiple languages within a single utterance, is a common linguistic phenomenon in multilingual societies [13]. Early studies on code-mixing focused primarily on sociolinguistic aspects, examining how and why speakers switch languages within conversations [11, 12, 13]. However, with the advent of digital communication, researchers have increasingly turned their attention to computational methods for processing and understanding code-mixed text [14].
@@
Several studies have explored various NLP tasks, such as part-of-speech tagging, language identification, and sentiment analysis, in code-mixed settings [15]. [16] provided one of the earliest comprehensive analyses of code-mixed text, highlighting the unique challenges it poses for traditional NLP pipelines, such as non-standard spelling, syntax variations, and the blending of multiple languages within a single text. More recent work by [17] introduced a code-mixed dataset, spanning multiple Indian languages, which has become a benchmark for evaluating NLP models in this domain.
@@
Information retrieval (IR) in code-mixed settings is relatively underexplored compared to other NLP tasks [18]. However, the need for effective IR systems that can handle multilingual and code-mixed queries has become increasingly important, particularly in the context of digital information exchange on social media platforms. [19] investigated the problem of query-focused summarization in code-mixed social media data, emphasizing the complexity of extracting relevant information from noisy, informal text. Work by [20] addressed code-mixed question answering, where the goal is to identify correct responses from a mixed-language corpus. Their approach involved using translation models to standardize the text before applying traditional IR techniques, demonstrating that even simple translation-based methods can significantly improve performance. However, these methods often fail to capture the nuances of code-mixed language, such as cultural context and colloquial expressions.
@@
Roman script transliteration of Indian languages, commonly referred to as "Romanagari" [21] for languages like Hindi, is a widespread practice in digital communication. Transliteration introduces additional challenges for NLP, as it often involves non-standard spellings and inconsistent usage. For instance, multiple transliterations may exist for the same word, depending on the speaker’s regional accent, literacy in the original script, or personal preference.
@@
Notable efforts in this area include the work by [22], which explored transliteration normalization for Hindi-English code-mixed text. They developed algorithms to map Romanized text back to its original script, enabling more accurate processing by traditional NLP models. However, normalization remains a challenging task due to the inherent variability in transliterated text. In the context of Bengali, the Roman script transliteration is less standardized than for Hindi, leading to even greater variability in spelling and grammar. [23] addressed this issue by creating a Roman Bengali dataset and proposed methods for transliteration normalization and language identification. Their work highlights the difficulties of processing Roman Bengali and the need for specialized approaches tailored to the characteristics of the language.
@@
Large Language Models (LLMs) [24, 25, 26] like GPT-3 have shown promise in various NLP tasks, including LI. Previous works have demonstrated the capability of GPT-3 in performing zero-shot and few-shot learning, making it a potentially powerful tool for LI in resource-constrained settings. However, the application of LLMs [27, 28, 29, 30] to code-mixed and morphologically rich languages remains underexplored. Recent studies, have started to explore the use of transformers and pre-trained models for multilingual LI, but the effectiveness of these models in Bengali languages requires further investigation.
@@ 3 Dataset
This shared task consists of a single dataset [31] for code mixed information retrieval. The corpus consists of 107900 documents in the training set and 20 queries in the training set. There are 30 queries in the testing set. The dataset is in roman transliterated bengali mixed with english language.
@@ 4 Task Definition
The task  [^1] https://cmir-iitbhu.github.io/cmir/ is to automatically determine the relevance of a query to a document within code-mixed data, specifically focusing on English and Roman transliterated Bengali.
@@ 5.1 Why Prompting?
  • Scalability and Customization: Prompting for information retrieval offers scalability and customization that traditional IR systems might lack. By designing prompts tailored to specific contexts or types of queries, LLMs can be dynamically adjusted to meet the needs of different retrieval tasks [36]. This customization is particularly useful in handling domain-specific language or code-mixed scenarios, where standard IR systems might require extensive re-training or re-configuration.
@@ 5.2 Merging Prompt and Mathematical Model-Based Approaches
We used the GPT-3.5 Turbo model via prompting through the OpenAI API [^2] https://platform.openai.com/docs/models/gpt-3-5-turbo to solve the document retrieval task. The process begins by first converting all the code-mixed sentences to english for both the queries and the documents. After this we try to determine the relevance scores, we used the following prompt:
@@
An example of the mathematical formulation and how it helps to detect relevant documents is shown in Table 1. The table reflects a range of documents to a code mixed query. The relevance scores help identify how closely each document addresses the query, while the probability scores provide insight into the potential usefulness of the documents based on the provided content. Overall, Documents 1-4 stands out as particularly relevant based on probability scores.
For the five results reported, we ran the GPT model at different temperature values namely 0.5, 0,6, 0.7, 0.8, and 0.9. The diagram for GPT-3.5 Turbo is shown in Figure 1. The figure representing the methodology is shown in Figure 2.
At lower temperatures, the model’s responses are more deterministic and focused. It generates outputs that are likely to be relevant and closely aligned with the input, making it useful for tasks requiring precision, such as retrieving specific information or handling queries with clear intent. At higher temperatures results in highly diverse and less predictable outputs. It can be useful in exploratory tasks where creativity and variation are needed, but it may also risk generating less coherent or relevant responses. In code-mixed scenarios, this could capture the full spectrum of linguistic creativity but might require careful handling to ensure relevance. So we used a temperature range of 0.5 to 0.9.
Query Document Relevance Score Probability Score
Kivabe bhalo bhabe ingreji shikhbo? Shudhumatro lekhar opor focus korle hobe na.
Ingreji songs shunle shikha sohoj hoy. 0.55 0.55 (1)
Kivabe bhalo bhabe ingreji shikhbo? Ingreji film dekhle vocabulary bere jaye.
Conversation practice korao khub helpful. 0.45 0.65 (1)
Kivabe bhalo bhabe ingreji shikhbo? Ingreji shikhar jonyo grammar book khub important.
Kintu speaking practice beshi dorkar. 0.35 0.55 (1)
Kivabe bhalo bhabe ingreji shikhbo? Bhasha shikhte bhalo tutor nirbachon kora joruri.
Bibhinno apps byabohar kore practice kora jaye. 0.45 0.65 (1)
Kivabe bhalo bhabe ingreji shikhbo? time lagbe. 0.20 0.20 (0)
Table 1: Example of Query and Document Relevance along with probability scores and relevance scores. Beside the probability score the relevance of a document to a query is stated in braces. 1 represents relevant and 0 represents not-relevant.
@@ 6 Results
Table 2 presents the evaluation metrics for different submissions [37] for the team named "TextTitans". The metrics used to assess the performance are MAP Score, ndcg Score, p@5 Score, and p@10 Score. Here’s what these results imply. MAP is a common metric in information retrieval that measures the precision of results across multiple queries. A higher MAP score indicates that relevant documents are consistently ranked higher across all queries. In the table, the MAP scores for the first four submissions are identical (0.701773), while the fifth submission slightly improves to 0.703734. This indicates that the fifth submission is marginally better in terms of ranking relevant results across multiple queries. The ndcg score measures the quality of the ranking based on the position of relevant documents. A higher ndcg score suggests that relevant documents are placed higher in the ranking. The scores are also very similar across submissions, with the first four submissions having an ndcg score of 0.797937, and the fifth submission showing a slight improvement to 0.799196. This suggests a minor improvement in ranking relevant documents for the fifth submission. p@5 measures how many of the top 5 ranked documents are relevant. A score of 1 would mean that all 5 of the top-ranked documents are relevant. All submissions have the same p@5 score of 0.793333, indicating that the top 5 results are equally accurate across all submissions. Precision@10 measures how many of the top 10 ranked documents are relevant. Like p@5, a higher score is better. Similar to p@5, all submissions have the same p@10 score of 0.766667, showing no variation in the top 10 results across the different submissions. The metrics are very consistent across all submissions, with only minor improvements in MAP and NDCG scores for the fifth submission. The fifth submission shows a slight improvement in ranking and retrieval performance, but the changes are minimal. The p@5 and p@10 scores indicate that the precision of the top 5 and top 10 results is identical across all submissions, suggesting that the models are performing similarly in identifying the most relevant documents. Overall, while there is a slight improvement in the last submission, the models generally perform similarly across all metrics.